ETEX(1) ETEX(1) NAME etex,einitex,evirtex−extendedTeX SYNOPSIS etex[options][commands] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. e-TEXisthefirstconcreteresultofaninternationalresearch&developmentproject,theNTS Project,whichwasestablishedundertheaegisofDANTEe.V.during1992.TheaimsoftheprojectaretoperpetuateanddevelopthespiritandphilosophyofTEX,whilstrespectingKnuth’s wishthatTEXshouldremainfrozen. e-TEXcanbeusedintwodifferentmodes:incompatibilitymodeitissupposedtobecompletely interchangablewithstandardTEX.Inextendedmodeseveralnewprimitivesareaddedthatfacilitate(amongotherthings)bidirectionaltypesetting. Anextendedmodeformatisgeneratedbyprefixingthenameofthesourcefilefortheformatwith anasterisk(*).Suchformatsareoftenprefixedwithan‘e’,henceetexastheextendedversionof texandelatexastheextendedversionoflatex.However,eplainisanexceptiontothisrule. Theeinitexandevirtexcommandsaree-TEX’sanaloguestotheinitexandvirtexcommands. Inthisinstallation,theyaresymlinkstotheetexexecutable. e-TEX’shandlingofitscommand-lineargumentsissimilartothatofTEX. OPTIONS Thisversionofe-TEXunderstandsthefollowingcommandlineoptions. --efmtformat Useformatasthenameoftheformattobeused,insteadofthenamebywhiche-TEX wascalledora%&line. --helpPrinthelpmessageandexit. --iniBeeinitex,fordumpingformats;thisisimplicitlytrueiftheprogramiscalledaseini-tex. --interactionmode Setstheinteractionmode.Themodecanbeoneofbatchmode,nonstopmode,scrollmode, anderrorstopmode.Themeaningofthesemodesisthesameasthatofthecorresponding \commands. --ipcSendDVIoutputtoasocketaswellastheusualoutputfile.Whetherthisoptionis availableisthechoiceoftheinstaller. --ipc-start As--ipc,andstartstheserverattheotherendaswell.Whetherthisoptionisavailable isthechoiceoftheinstaller. --kpathsea-debugbitmask Setspathsearchingdebuggingflagsaccordingtothebitmask.SeetheKpathseamanual fordetails. --maketexfmt Enablemktexfmt,wherefmtmustbeoneoftexortfm. --mltex EnableMLTEXextensions. --no-maketexfmt Disablemktexfmt,wherefmtmustbeoneoftexortfm. Web2C7.32x 29March1999 1 ETEX(1) ETEX(1) --output-commentstring UsestringfortheDVIfilecommentinsteadofthedate. --prognamename Pretendtobeprogramname.Thisaffectsboththeformatusedandthesearchpaths. --shell-escape Enablethe\write18{command}construct.ThecommandcanbeanyBourneshellcommand.Thisconstructisnormallydisallowedforsecurityreasons. --translate-filetcxname Usethetcxnametranslationtable. --version Printversioninformationandexit. ENVIRONMENT SeetheKpathsearchlibrarydocumentation(the‘Pathspecifications’node)forprecisedetailsof howtheenvironmentvariablesareused.Thekpsewhichutilitycanbeusedtoquerythevalues ofthevariables. Onecaveat:Inmoste-TEXformats,youcannotuse˜ inafilenameyougivedirectlytoe-TEX, because˜ isanactivecharacter,andhenceisexpanded,nottakenaspartofthefilename.Other programs,suchasMETAFONT,donothavethisproblem. TEXMFOUTPUT Normally,e-TEXputsitsoutputfilesinthecurrentdirectory.Ifanyoutputfilecannotbe openedthere,ittriestoopenitinthedirectoryspecifiedintheenvironmentvariable TEXMFOUTPUT.Thereisnodefaultvalueforthatvariable.Forexample,ifyousay texpaperandthecurrentdirectoryisnotwritable,ifTEXMFOUTPUThasthevalue /tmp,e-TEXattemptstocreate/tmp/paper.log(and/tmp/paper.dvi,ifanyoutputispro-duced.) TEXINPUTS Searchpathfor\inputand\openinfiles.Thisshouldprobablystartwith‘‘.’’,sothat userfilesarefoundbeforesystemfiles.Anemptypathcomponentwillbereplacedwith thepathsdefinedinthetexmf.cnffile.Forexample,setTEXINPUTSto ".:/home/usr/tex:"toprependthecurrentdirecoryand‘‘/home/user/tex’’tothestandardsearchpath. TEXFONTS Searchpathforfontmetric(.tfm)files. TEXFORMATS Searchpathforformatfiles. TEXPOOL searchpathforeinitexinternalstrings. TEXEDIT Commandtemplateforswitchingtoeditor.Thedefault,usuallyvi,issetwhene-TEXis compiled. FILES Thelocationofthefilesmentionedbelowvariesfromsystemtosystem.Usethekpsewhichutilitytofindtheirlocations. etex.pool Encodedtextofe-TEX’smessages. texfonts.map Filenamemappingdefinitions. Web2C7.32x 29March1999 2 ETEX(1) ETEX(1) *.tfmMetricfilesfore-TEX’sfonts. *.efmtPredigestede-TEXformat(.efmt)files. BUGS Thisversionofe-TEXfailstotraparithmeticoverflowwhendimensionsareaddedorsubtracted. Caseswherethisoccursarerare,butwhenitdoesthegeneratedDVIfilewillbeinvalid. SEEALSO tex(1),mf(1), Web2C7.32x 29March1999 3